A body freely falling frim the rest has a velocity `v` after it falls through a height `h` the distance it has to fall down for its velocity becomes double, is.......

Grade:11

1 Answers

Samyak
6 years ago
Body starts from rest thusu=0 after moving h height it has a velocity v.Therefore using v^2=u^2+2axWe can write v^2=2ghThis gives h= v^2/2g Now when velocity becomes double the distance it travels from initial point is given by(2v)^2=2gr (r is distance in this case)r=4v^2/2gFinally the distance it needs to travel after travelling h is r-h. Which equals 4v^2/2g -v^2/2gAfter simplifying which becomes 3v^2/2g